# Tensor Attention Training: Provably Efficient Learning of Higher-order Transformers

### v2

@@ D.2 Definitions for loss function
For all $j_{0}\in[n]$, for all $i_{0}\in[d]$. We define $\mathsf{Loss}(x)_{j_{0},i_{0}}$ to be $:=0.5\mathsf{V}(x)_{j_{0},i_{0}}^{2}$.
@@ D.3 Further information on gradient computation
To further simplify the writing of proofs, we represent $(x)$ as ()($\cdot$)( ⋅ ).
@@
To further simplify the writing of proofs, we represent $(x)$ as ()($\cdot$)( ⋅ ).
@@
To further simplify the writing of proofs, we represent $(x)$ as ()($\cdot$)( ⋅ ).
@@
To further simplify the writing of proofs, we represent $(x)$ as ()($\cdot$)( ⋅ ).
@@ F.6 Gradient computation in almost linear time by low rank tensor approximation
Theorem F.8 (Main result for fast gradient computation, Restatement of Theorem 5.2).
@@
Recall that $\widetilde{g}\in\mathbb{R}^{d\times d^{2}}$ and $\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d^{2}}$
@@ G.1 Tools for backward complexity
Suppose that a fixed matrix $H\in\mathbb{R}^{n\times n^{2}}$ with entries in the interval $[1,B_{a}]$ satisfying that more than half entries of $H$ in each row are equal to $B_{a}$. Let a matrix $V\in\mathbb{R}^{n^{2}\times d}$ with entries in $\{0,1\}$. For $\lambda\in\mathbb{R}$, let us define $M_{\lambda}:=\exp(\lambda H)\in\mathbb{R}^{n\times n^{2}}$. We denote the function $f:\mathbb{R}\to\mathbb{R}$ as
@@
As at least half of the entries in each row of $H$ are equal to $B_{a}$ and all entries lie within the interval $[1,B_{a}]$, we can bound:
@@ G.2 Main result for lower bound
Theorem G.3 (Main result for hardness, Restatement of Theorem 6.3).
@@
Suppose that $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ is an identity tensor. Also suppose that the input matrices to Lemma 6.2 are $Q,K_{1},K_{2},V_{1},V_{2}$. And we set $A_{1}=Q$, $A_{2}=K_{1}$,$A_{3}=K_{2}$, $A_{4}=V_{1}$,$A_{5}=V_{2}$, $Y=I$, and $X=\lambda\cdot\underbrace{\mathrm{mat}(\mathsf{I}_{d})}_{d\times d^{2}}$, with some $\lambda\in[0,1]$. Let $f:[0,1]\to\mathbb{R}$ be defined in Lemma G.1 where $H$ is the matrix $A_{1}(A_{2}\oslash A_{3})^{\top}$, so that $M_{\lambda}$ is the matrix $\exp(A_{1}X(A_{2}\otimes A_{3})^{\top})$ by Fact C.19. It follows from Lemma G.1 and $d=\Theta(\log n)$ that

### v3

@@ D.2 Definitions for loss function
For all $j_{0}\in[n]$, for all $i_{0}\in[d]$. We define $\mathsf{Loss}(x)_{j_{0},i_{0}}$ to be $0.5\mathsf{V}(x)_{j_{0},i_{0}}^{2}$.
@@ D.3 Further information on gradient computation
To further simplify the writing of proofs, we represent $(x)$ as $(\cdot)$.
@@
To further simplify the writing of proofs, we represent $(x)$ as $(\cdot)$.
@@
To further simplify the writing of proofs, we represent $(x)$ as $(\cdot)$.
@@
To further simplify the writing of proofs, we represent $(x)$ as $(\cdot)$.
@@ F.6 Gradient computation in almost linear time by low rank tensor approximation
Theorem F.8 (Main result for fast gradient computation, formal version of Theorem 4.2).
@@
Recall that $\widetilde{g}\in\mathbb{R}^{d\times d^{2}}$ and $\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d^{2}}$ .
@@ G.1 Tools for backward complexity
Suppose that a fixed matrix $H\in\mathbb{R}^{n\times n^{2}}$ with entries in the interval $1,B_{a}]$ satisfying that more than half entries of $H$ in each row are equal to $B_{a}$. Let a matrix $V\in\mathbb{R}^{n^{2}\times d}$ with entries in $\{0,1\}$. For $\lambda\in\mathbb{R}$, let us define $M_{\lambda}:=\exp(\lambda H)\in\mathbb{R}^{n\times n^{2}}$. We denote the function $f:\mathbb{R}\to\mathbb{R}$ as
@@
As at least half of the entries in each row of $H$ are equal to $B_{a}$ and all entries lie within the interval $1,B_{a}]$, we can bound:
@@ G.2 Main result for lower bound
Theorem G.3 (Main result for hardness, formal of Theorem 5.3).
@@
Suppose that $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ is an identity tensor. Also suppose that the input matrices to Lemma 5.2 are $Q,K_{1},K_{2},V_{1},V_{2}$. And we set $A_{1}=Q$, $A_{2}=K_{1}$$A_{3}=K_{2}$, $A_{4}=V_{1}$$A_{5}=V_{2}$, $Y=I$, and $X=\lambda\cdot\underbrace{\mathrm{mat}(\mathsf{I}_{d})}_{d\times d^{2}}$, with some $\lambda\in[0,1]$. Let $f:[0,1]\to\mathbb{R}$ be defined in Lemma G.1 where $H$ is the matrix $A_{1}(A_{2}\oslash A_{3})^{\top}$, so that $M_{\lambda}$ is the matrix $\exp(A_{1}X(A_{2}\otimes A_{3})^{\top})$ by Fact C.19. It follows from Lemma G.1 and $d=\Theta(\log n)$ that